# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has recently emerged as a powerful alternative to Neural Radiance Fields (NeRF) for 3D scene representation, offering high-fidelity photorealistic rendering with real-time performance. Beyond novel view synthesis, the explicit and compact nature of 3DGS enables a wide range of downstream applications that require geometric and semantic understanding. This survey provides a comprehensive overview of recent progress in 3DGS applications. It first introduces 2D foundation models that support semantic understanding and control in 3DGS applications, followed by a review of NeRF-based methods that inform their 3DGS counterparts. We then categorize 3DGS applications into segmentation, editing, generation, and other functional tasks. For each, we summarize representative methods, supervision strategies, and learning paradigms, highlighting shared design principles and emerging trends. Commonly used datasets and evaluation protocols are also summarized, along with comparative analyses of recent methods across public benchmarks. To support ongoing research and development, a continually updated repository of papers, code, and resources is maintained at https://github.com/heshuting555/Awesome-3DGS-Applications.
@@ 1 Introduction
3D Gaussian Splatting (3DGS)[1] has recently emerged as a powerful paradigm for real-time neural rendering, achieving high-fidelity photorealistic synthesis with superior efficiency. Unlike implicit field-based methods such as Neural Radiance Fields (NeRFs)[2], which rely on volumetric integration over coordinate-based neural networks, 3DGS represents scenes as explicit collections of anisotropic 3D Gaussians and renders them via differentiable rasterization. This explicit yet learnable formulation enables efficient optimization and fast inference, while preserving fine-grained geometric and appearance details. As a next-generation neural representation, 3DGS has demonstrated remarkable potential across a broad spectrum of applications, including virtual and augmented reality, robotics, autonomous navigation, and urban mapping.
@@
Although several recent surveys [10, 11, 12, 13] have documented the rapid development of 3DGS, they primarily focus on global taxonomies, real-time rendering pipelines, or compression strategies, while offering limited insights into 3DGS-driven downstream applications. A few works [14, 15, 16] attempt to cover application domains, but typically lack systematic analysis of underlying design principles, methodological innovations, or benchmarking protocols. To address this gap, we present the first dedicated survey that systematically reviews downstream applications of 3DGS beyond classical view synthesis. Specifically, we focus on three rapidly evolving directions, namely segmentation, editing, and generation, as illustrated in Fig. 1.
@@
Figure 1: Overview of the three main 3DGS applications. (a) Segmentation: Segment each semantic region. (b) Editing: Perform editing via text or image prompt. (c) Generation: Support text-to-3D and image-to-3D generation.
Figure 2: Overview of the structure of this survey.
@@
$\bullet$ Contribution. This survey provides a comprehensive and structured review of recent literature on 3DGS with a focus on its emerging downstream applications. We begin by introducing 2D foundation models that enable semantic control and understanding in 3DGS-based systems, and briefly revisit relevant NeRF-based methods to establish conceptual continuity. The core of our survey is organized around three major task categories: segmentation, editing, and generation. Within each category, we systematically compare representative works in terms of technical designs, learning paradigms, and supervision strategies. We further analyze benchmark settings and performance results in each domain, aiming to provide a self-contained and accessible resource for both newcomers and experienced researchers. Finally, we discuss open challenges and future research opportunities, with the goal of fostering deeper exploration and broader adoption of 3DGS across high-level vision and graphics application tasks.
@@
$\bullet$ Organization. Fig. 2 provides an overview of the structure of this survey. In Sec. 2, we introduce essential background knowledge, including key definitions, commonly used 2D foundation models, and relevant NeRF-based research that informs 3DGS development. Sec. 3 presents a task-centric review of recent 3DGS-based methods, with a focus on segmentation, editing, and generation. In Sec. 4, we summarize evaluation protocols and benchmark settings, and provides quantitative comparisons across tasks and datasets to assess the performance of existing approaches. Sec. 5 discusses open challenges and outlines potential future research directions. Finally, Sec. 6 concludes the survey with a summary of key insights and takeaways.
@@ 2.1.1 Application Tasks Category
Depending on the nature of the output space $\mathcal{Y}$, downstream applications built upon 3DGS can be broadly categorized into three representative directions: segmentation, editing, and generation.
@@ 2.3 Related Research Areas
$\bullet$ Generation Based on NeRF. NeRF-based generative models have rapidly advanced 3D-aware content synthesis. Early works utilize implicit volumetric representations with adversarial training to achieve 3D consistent image generation from unposed views, improving shape–appearance disentanglement and scalability [76]. To boost efficiency and detail, later methods constrain sampling to 2D manifolds via ray-surface intersections [77]. A major shift occurs with Score Distillation Sampling (SDS), which enables NeRF optimization from text prompts using 2D diffusion models, eliminating the need for 3D supervision [78]. Follow-up work adopts coarse-to-fine strategies and hybrid representations to enhance quality and speed [79]. Latent-NeRF [80] uses Sketch-Shape to define the coarse structure of abstract geometric objects, enabling text- and shape-guided control over the generation process. Wang et al. [81] apply the chain rule to the learned gradients and backpropagate the diffusion model’s score through the Jacobian of a differentiable renderer, instantiated as a voxel radiance field. In parallel, GAN-based models introduce tri-plane features and pose-aware training to generate high-resolution, multi-view consistent outputs [82]. Recent diffusion-guided approaches further extend to full-scene generation by combining semantic priors and progressive inpainting for photorealistic results [83].
3 3DGS Application Tasks
@@
3DGS segmentation encompasses a broad spectrum of task definitions, as discussed in Sec. 2.1.1. Given the overlapping nature of segmentation tasks and that a single method may apply to multiple scenarios, we organize existing approaches based on their technical characteristics rather than task formulations.
@@ 3.1.1 Feature Distillation-based Methods
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [88] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [89] improves LangSplat by extracting dense pixel-level semantics from the entire image, followed by SAM-based mask pooling to preserve contextual cues. It also jointly optimizes language Gaussians on object surfaces using geometry supervision and contrastive loss, enhancing object’s spatial consistency in 3D. seconGS [90] further refines LangSplat by using SAM2-generated masklets to provide semantically consistent supervision. To overcome the limitations of 2D querying in 3D space, it introduces a two-step querying mechanism that first retrieves distilled ground-truth and then queries individual Gaussians for spatially accurate 3D understanding. VLGaussian [28] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [91], inspired by superpoint representations, clusters Gaussians into Super-Gaussians using spatial and instance cues based on Scaffold-GS [3]. High-dimensional language features are assigned to each Super-Gaussian to support structured and comprehensive scene understanding while preserving linguistic granularity. FreeGS [92] introduces a semantic-embedded 3DGS framework by coupling semantic features with instance identities via the proposed IDSF field. A two-step alternating optimization and a 2D–3D contrastive loss ensure view-consistent semantics without requiring 2D labels. Beyond the aforementioned approaches that solely introduce semantic information, GLS [93] incorporates geometric priors (surface normals) to effectively refine the rendered depth.
@@
$\bullet$ Efforts toward Efficiency and Compactness. To mitigate the substantial memory overhead incurred by increasing feature channels in raw semantic features, LEGaussian [84] quantizes dense language features into a discrete feature space and semantic indices, and applies a smoothing loss to retain rendering quality with quantized representation. CLIP-GS [94] introduces semantic attribute compactness to efficiently encode scene semantics with Gaussians, enabling fast training and inference. It further improves 3D segmentation by employing a 3D-coherent self-training strategy that enhances cross-view semantic consistency. Later, FMGS [95] integrates 3D Gaussian scene representation with multi-resolution hash encodings (MHE) to enable efficient semantic embedding. Unlike previous methods, it avoids scene-specific quantization or autodecoders, thereby preserving the semantic fidelity of foundation model features. GOI [96] introduces a trainable feature clustering codebook that compresses high-dimensional, noisy semantic features into compact representations guided by scene priors. It effectively preserves semantic expressiveness while significantly reducing storage and rendering overhead. FastLGS [97] constructs a semantic feature grid that stores multi-view CLIP features extracted from SAM-generated masks. These grid features are then projected into a low-dimensional space and used to supervise semantic field training. Building upon FastLGS, FMLGS [98] enables part-level open-vocabulary queries in 3DGS by constructing consistent object-part semantics via SAM2, and introduces a semantic deviation strategy to enrich fine-grained features, allowing natural language queries over both objects and their parts. To address the high memory cost of jointly encoding color and semantics in existing 3DGS-based methods[27, 84, 94], DF-3DGS [99] decouples color and semantic fields to reduce Gaussian usage for semantics, and introduces a hierarchical compression pipeline combining dynamic quantization and scene-specific autoencoding for efficient representation. In summary, knowledge distillation provides a practical mechanism to bridge 2D foundation models and 3D Gaussian representations. While early methods focus on direct supervision, recent advances have introduced architectural, supervisory, and efficiency-oriented innovations to improve precision and scalability. These works collectively advance open-vocabulary 3D understanding under constrained computational and data resources.
@@
$\bullet$ Consistency-Aware End-to-End Learning. To reduce reliance on heuristic pre- and post-processing, recent works [30, 104, 24, 32, 105, 17, 106] propose end-to-end learning frameworks that directly construct consistent and distinguishable 3D feature fields. SAGA [30] distills 2D masks into 3D features that introduce a scale-aware contrastive training strategy that distills SAM’s segmentation capability into a scale-gated affinity representation, which handles multi-granularity ambiguity. SA3D [104] introduces an iterative refinement strategy that alternates between inverse rendering of 2D masks and cross-view self-prompting, enabling progressively more consistent and accurate object segmentation across multiple views. Click-Gaussian [24] enables efficient interactive segmentation of 3D Gaussians by incorporating multi-granularity feature fields derived from 2D masks. It further introduces Global Feature-guided Learning (GFL) to alleviate inconsistencies across different views. Unified-Lift [20] develops an object-aware lifting pipeline that eliminates the need for additional alignment steps. It formulates a codebook-based object representation and aligns object-level semantics with Gaussian features via contrastive learning, leading to improved 3D instance understanding. OpenGaussian [32] focuses on 3D open-vocabulary segmentation by leveraging SAM-predicted masks to supervise 3D instance features. It adopts a coarse-to-fine feature discretization strategy via a two-stage codebook and introduces an instance-level 3D–2D association module that bridges Gaussian points with 2D masks and CLIP embeddings. econSG [107] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [108] aligns SAM-generated masks across views via a zero-shot tracker, extracts semantics with CLIP, and applies contrastive codebook learning to enforce cross-view consistency by promoting intra-class compactness and inter-class separation. VoteSplat [109] integrates Hough voting with 3DGS by learning per-Gaussian 3D offsets whose projections align with SAM-derived 2D vote maps and are depth-regularized, so that votes aggregate near object centroids. ILGS [110] mitigates view-inconsistent language embeddings by enforcing cross-view semantic alignment with an identity-aware semantic consistency loss and refining boundaries via progressive mask expansion.
@@ 3.1.3 Feed-Forward-based Methods
To overcome the inefficiency of per-scene optimization, especially under dense calibrated images conditions, several methods [117, 118, 119, 120, 121, 122] adopt feed-forward architectures to enable fast and generalizable 3D semantic field construction. LSM [123] is the first feed-forward framework that integrates DUSt3R [124], point transformer [125], and LSeg [87] for language-driven scene understanding. SLGaussian [117] builds a feed-forward pipeline that infers 3D semantic Gaussians directly from sparse viewpoints based on MVSplat [4]. Dr. Splat [119] departs from traditional rendering-based supervision by directly associating CLIP embeddings with 3D Gaussians. It introduces a language feature registration mechanism that assigns CLIP features to dominant Gaussians intersected by pixel rays. Additionally, it employs Product Quantization (PQ) learned from large-scale image datasets to compress features. LangScene-X [121] first uses a TriMap video diffusion network based on CogVideoX [126] to densify sparse inputs, synthesizing appearance, geometry, and semantics. It then employs a Language-Quantized Compressor (LQC) to encode language embeddings on large-scale datasets. SceneSplat [127] introduces a self-supervised framework that facilitates rich 3D feature learning from unlabeled scenes. To support this effort, SceneSplat-7K is introduced as the first large-scale dataset specifically designed for 3DGS in indoor environments.

### v3

@@ Abstract
In the context of novel view synthesis, 3D Gaussian Splatting (3DGS) has recently emerged as an efficient and competitive counterpart to Neural Radiance Field (NeRF), enabling high-fidelity photorealistic rendering in real time. Beyond novel view synthesis, the explicit and compact nature of 3DGS enables a wide range of downstream applications that require geometric and semantic understanding. This survey provides a comprehensive overview of recent progress in 3DGS applications. It first introduces 2D foundation models that support semantic understanding and control in 3DGS applications, followed by a review of NeRF-based methods that inform their 3DGS counterparts. We then categorize 3DGS applications into three foundational tasks: segmentation, editing, and generation, alongside additional functional applications built upon or tightly coupled with these foundational capabilities. For each, we summarize representative methods, supervision strategies, and learning paradigms, highlighting shared design principles and emerging trends. Commonly used datasets and evaluation protocols are also summarized, along with comparative analyses of recent methods across public benchmarks. To support ongoing research and development, a continually updated repository of papers, code, and resources is maintained at https://github.com/heshuting555/Awesome-3DGS-Applications.
@@ 1 Introduction
3D Gaussian Splatting (3DGS)[1] has recently emerged as a powerful paradigm for real-time neural rendering, achieving high-fidelity photorealistic synthesis with superior efficiency. Neural Radiance Field (NeRF) [2] represents a scene as an implicit radiance field parameterized by a neural network that maps 3D coordinates and viewing directions to density and view-dependent radiance, which is rendered via volume rendering. In contrast, 3DGS represents a scene as an explicit set of anisotropic 3D Gaussians and renders them via differentiable rasterization. This explicit yet learnable formulation enables efficient optimization and fast inference, while preserving fine-grained geometric and appearance details. As a next-generation explicit 3D representation, 3DGS has demonstrated remarkable potential across a broad spectrum of applications, including virtual and augmented reality, robotics, autonomous navigation, and urban mapping.
@@
Although several recent surveys [10, 11, 12, 13] have documented the rapid development of 3DGS, they primarily focus on global taxonomies, real-time rendering pipelines, or compression strategies, while offering limited insights into 3DGS-driven downstream applications. A few works [14, 15, 16] attempt to cover application domains, but typically lack systematic analysis of underlying design principles, methodological innovations, or benchmarking protocols. To address this gap, we present the first dedicated survey that systematically reviews downstream applications of 3DGS beyond classical view synthesis. Furthermore, we organize 3DGS application tasks into foundational and functional categories. The former represents general capabilities that are broadly used in various application scenarios, while the latter comprises scenario-driven pipelines tailored to specific applications built upon or tightly coupled with these capabilities. Accordingly, we focus primarily on three foundational tasks, as summarized in Fig. 1, and discuss functional tasks in the Appendix.
@@
Figure 1: Overview of the three foundational 3DGS application tasks. (a) Segmentation: Segment each semantic region. (b) Editing: Perform editing via text or image prompt. (c) Generation: Support text-to-3D and image-to-3D generation.
§1§2§3§3.1§3.2§3.3§4§5§6§A§2.1§2.2§2.3§3.1.1§3.1.2§3.1.3§3.1.4§3.2.1§3.2.2§3.2.3§3.3.1§3.3.2§A.1§A.2
Figure 2: Overview of the survey structure.
@@
$\bullet$ Contribution. This survey provides a comprehensive and structured review of recent literature on 3DGS with a focus on its emerging downstream applications. We begin by introducing 2D foundation models that enable semantic control and understanding in 3DGS-based systems, and briefly revisit relevant NeRF-based methods to establish conceptual continuity. The core of our survey is organized around three foundational task categories: segmentation, editing, and generation. Within each category, we systematically compare representative works in terms of technical designs, learning paradigms, and supervision strategies. We further analyze benchmark settings and performance results in each domain, aiming to provide a self-contained and accessible resource for both newcomers and experienced researchers. Finally, we discuss open challenges and future research opportunities, with the goal of fostering deeper exploration and broader adoption of 3DGS across high-level vision and graphics application tasks.
@@
Fig. 2 provides an overview of the structure of this survey. In Sec. 2, we introduce essential background knowledge, including key definitions, commonly used 2D foundation models, and relevant NeRF-based research that informs 3DGS development. Sec. 3 presents a task-centric review of recent 3DGS-based methods, with a focus on segmentation, editing, and generation. In Sec. 4, we summarize evaluation protocols and benchmark settings, and provide quantitative comparisons across tasks and datasets to assess the performance of existing approaches. Sec. 5 discusses open challenges and outlines potential future research directions. Sec. 6 concludes the survey with a summary of key insights and takeaways. Finally, we present a discussion of functional applications and complete dataset statistics tables in the Appendix.
@@ 2.1.1 Application Tasks Category
Downstream applications built upon 3DGS can be broadly categorized into foundational and functional tasks. In this survey, we focus on three representative foundational tasks: segmentation, editing, and generation, which differ in their output space $\bm{\mathcal{Y}}$.
@@ 2.3 Related Research Areas
Figure 3: A comparison of three 3DGS segmentation pipelines.
$\bullet$ Generation Based on NeRF. NeRF-based generative models have rapidly advanced 3D-aware content synthesis. Early works utilize implicit volumetric representations with adversarial training to achieve 3D consistent image generation from unposed views, improving shape–appearance disentanglement and scalability [76]. To boost efficiency and detail, later methods constrain sampling to 2D manifolds via ray-surface intersections [77]. A major shift occurs with Score Distillation Sampling (SDS), which enables NeRF optimization from text prompts using 2D diffusion models, eliminating the need for 3D supervision [78]. Follow-up work adopts coarse-to-fine strategies and hybrid representations to enhance quality and speed [79]. Latent-NeRF [80] uses Sketch-Shape to define the coarse structure of abstract geometric objects, enabling text- and shape-guided control over the generation process. Wang et al. [81] apply the chain rule to the learned gradients and backpropagate the diffusion model’s score through the Jacobian of a differentiable renderer, instantiated as a voxel radiance field. In parallel, GAN-based models introduce tri-plane features and pose-aware training to generate high-resolution, multi-view consistent outputs [82]. Recent diffusion-guided approaches further extend to full-scene generation by combining semantic priors [83].
3 Foundational Application Tasks
@@
In 3DGS-based scene understanding, segmentation and object detection are closely related tasks and often share similar representations and learning paradigms. We begin with segmentation, as it has been more extensively studied and provides a natural basis for organizing related scene understanding techniques. For completeness, we then review object detection in Sec. 3.1.4 as a closely related task to segmentation, emphasizing both shared components and task-specific differences. Segmentation encompasses a broad spectrum of task definitions, as discussed in Sec. 2.1.1. Given the substantial overlap among task settings, where a single method may apply to multiple scenarios, we organize existing approaches by their technical characteristics rather than by task formulations. Representative segmentation pipelines are provided in Fig. 3.
@@ 3.1.1 Feature Distillation-based Methods
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [88] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [89] derives dense pixel-level semantics with SAM pooling, and jointly optimizes surface language Gaussians with geometric supervision and contrastive loss for improved 3D spatial consistency. seconGS [90] refines LangSplat with SAM2 masklet supervision and a two-step querying scheme that first retrieves distilled ground-truth and then queries individual Gaussians for more accurate 3D understanding. VLGaussian [28] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [91] clusters Gaussians into Super-Gaussians with spatial and instance cues built on Scaffold-GS [3] and assigns language features per cluster for structured scene understanding. FreeGS [92] introduces a semantic-embedded 3DGS framework by coupling semantic features with instance identities via the proposed IDSF field. A two-step alternating optimization and a 2D–3D contrastive loss ensure view-consistent semantics without requiring 2D labels. econSG [93] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [94] aligns SAM-generated masks across views via a zero-shot tracker, extracts semantics with CLIP, and applies contrastive codebook learning to enforce cross-view consistency by promoting intra-class compactness and inter-class separation. ILGS [95] mitigates view-inconsistent language embeddings by enforcing cross-view semantic alignment with an identity-aware semantic consistency loss and refining boundaries via progressive mask expansion. COS3D [96] learns collaborative instance and language fields, using SAM masks with contrastive instance supervision and an instance-to-language mapping. At inference, it uses adaptive language-to-instance prompt refinement to produce accurate output. Beyond the aforementioned approaches that solely introduce semantic information, GLS [97] incorporates geometric priors (surface normals) to effectively refine the rendered depth.
@@
$\bullet$ Efforts toward Efficiency and Compactness. To mitigate the substantial memory overhead incurred by increasing feature channels in raw semantic features, LEGaussian [84] quantizes dense language features into a discrete feature space and semantic indices, and applies a smoothing loss to retain rendering quality with quantized representation. CLIP-GS [98] introduces semantic attribute compactness to efficiently encode scene semantics with Gaussians, enabling fast training and inference. It further improves 3D segmentation by employing a 3D-coherent self-training strategy that enhances cross-view semantic consistency. Later, FMGS [99] integrates 3D Gaussian scene representation with multi-resolution hash encodings (MHE) to enable efficient semantic embedding. Unlike previous methods, it avoids scene-specific quantization or autodecoders, thereby preserving the semantic fidelity of foundation model features. GOI [100] learns a clustering codebook to compress noisy high-dimensional semantic features into compact, scene-prior-guided representations, reducing storage and rendering cost. FastLGS [101] constructs a semantic feature grid that stores multi-view CLIP features extracted from SAM-generated masks. These grid features are then projected into a low-dimensional space and used to supervise semantic field training. Building upon FastLGS, FMLGS [102] enables part-level open-vocabulary queries in 3DGS by constructing consistent object-part semantics via SAM2, and introduces a semantic deviation strategy to enrich fine-grained features, allowing natural language queries over both objects and their parts. To address the high memory cost of jointly encoding color and semantics in existing 3DGS-based methods[27, 84, 98], DF-3DGS [103] decouples color and semantic fields to reduce Gaussian usage for semantics, and introduces a hierarchical compression pipeline combining dynamic quantization and scene-specific autoencoding for efficient representation. LansSplatV2 [104] represents each Gaussian’s CLIP feature using a global codebook with sparse top-$K$ coefficients, rendering only the coefficients for efficient open-vocabulary querying. In summary, knowledge distillation provides a practical mechanism to bridge 2D foundation models and 3D Gaussian representations. While early methods focus on direct supervision, recent advances introduce architectural, supervisory, and efficiency-oriented innovations to improve precision and scalability. These works collectively advance open-vocabulary 3D understanding under constrained computational and data resources.
@@
$\bullet$ Consistency-Aware End-to-End Learning. To reduce reliance on heuristic pre- and post-processing, recent works [30, 109, 24, 32, 110, 17, 111] propose end-to-end learning frameworks that directly construct consistent and distinguishable 3D feature fields. SAGA [30] distills 2D masks into 3D features that introduce a scale-aware contrastive training strategy that distills SAM’s segmentation capability into a scale-gated affinity representation, which handles multi-granularity ambiguity. SA3D [109] introduces an iterative refinement strategy that alternates between inverse rendering of 2D masks and cross-view self-prompting, enabling progressively more consistent and accurate object segmentation across multiple views. Click-Gaussian [24] enables efficient interactive segmentation of 3D Gaussians by incorporating multi-granularity feature fields derived from 2D masks. It further introduces Global Feature-guided Learning (GFL) to alleviate inconsistencies across different views. Unified-Lift [20] develops an object-aware lifting pipeline that eliminates the need for additional alignment steps. It formulates a codebook-based object representation and aligns object-level semantics with Gaussian features via contrastive learning, leading to improved 3D instance understanding. OpenGaussian [32] focuses on 3D open-vocabulary segmentation by leveraging SAM-predicted masks to supervise 3D instance features. It adopts a coarse-to-fine feature discretization strategy via a two-stage codebook and introduces an instance-level 3D–2D association module that bridges Gaussian points with 2D masks and CLIP embeddings. VoteSplat [112] integrates Hough voting with 3DGS by learning per‑Gaussian 3D offsets whose projections align with SAM‑derived 2D vote maps and are depth‑regularized.
Figure 4: Examples of different editing operations in 3DGS, with examples adapted from GaussianEditor [8], PortraitGen [113], SGSST [114], GScream [115], and DYG [116].
@@ 3.1.3 Feed-Forward-based Methods
To overcome the inefficiency of per-scene optimization, especially under dense calibrated images conditions, several methods [123, 124, 125, 126, 127, 128] adopt feed-forward architectures to enable fast and generalizable 3D semantic field construction. LSM [129] is the first feed-forward framework that integrates DUSt3R [130], point transformer [131], and LSeg [87] for language-driven scene understanding. Following LSM’s paradigm, SIU3R [132] is proposed to learn scene understanding and 3D reconstruction simultaneously without feature alignment. SLGaussian [123] builds a feed-forward pipeline that infers 3D semantic Gaussians directly from sparse viewpoints based on MVSplat [4]. Dr. Splat [125] departs from traditional rendering-based supervision by directly associating CLIP embeddings with 3D Gaussians. It introduces a language feature registration mechanism that assigns CLIP features to dominant Gaussians intersected by pixel rays. Additionally, it employs Product Quantization (PQ) learned from large-scale image datasets to compress features. LangScene-X [127] first uses a TriMap video diffusion network based on CogVideoX [133] to densify sparse inputs, synthesizing appearance, geometry, and semantics. It then employs a Language‑Quantized Compressor (LQC) to encode language embeddings on large-scale datasets. SceneSplat [134] introduces a self-supervised framework that facilitates rich 3D feature learning from unlabeled scenes. To support this effort, SceneSplat-7K is introduced as the first large-scale dataset specifically designed for 3DGS in indoor environments.
3.1.4 Related Task: 3DGS Object Detection
3D object detection focuses on identifying and localizing objects in 3D space. Similar to segmentation, detection methods can be categorized into prompt-based and generic approaches. Prompt-based detection, often referred to as open-vocabulary detection or visual grounding, takes textual inputs (e.g., category names or language expressions) and outputs object locations. Most 3DGS-based segmentation methods also support this type of detection by generating the maximum response point, and are thus not detailed here. Specifically, SpatialReasoner [135] combines LLM-driven spatial reasoning with a hierarchical feature field enhanced by visual properties. It distills CLIP and SAM features and leverages fine-tuned LLMs to infer spatial instructions, enabling precise localization of target instances based on relational language cues. ReasonGrounder [136] leverages large vision-language models (LVLM) to interpret both explicit and implicit instructions. It localizes objects, even when partially or fully occluded, via hierarchical 3D feature fields that enable scale-aware Gaussian grouping. GVR [137] is a zero-shot 3D visual grounding framework that reformulates 3DVG as a 2D retrieval task. Relying on existing 2D visual foundation models and a built knowledge book, it removes the need for 3D annotations and per-scene training. For generic object detection, representative methods include: Gaussian-Det [138], which models objects continuously using input Gaussians as feature descriptors across partial surfaces; 3DGS-DET [139], which incorporates boundary guidance and box-focused sampling to refine Gaussian distributions and suppress background noise; and MATT-GS [140], which employs masked attention 3D Gaussian Splatting to enhance localization.